If c=R20 billion, I=R30 billion, g=R25 billion, x=R15 billion, z=R12 billion, c=0,75 the value of the multiplier is?


1
Expert's answer
2021-08-05T13:07:35-0400

Solution:

The value of multiplier = 4

The multiplier formula = "\frac{1}{1 - MPC}"

MPC = 0.75

The value of multiplier = "\frac{1}{1 - 0.75} = \frac{1}{0.25} = 4" 1/(1 – 0.75) = 1/0.25 = 4


The value of multiplier = 4